# YITP-24-61

### v1

@@ Abstract
where $\rho$ denotes the radial coordinate in the 4-dimensional Euclidean space. The second condition in (3) guarantees the regularity of the solution at the center of the true-vacuum bubble. We call Euclidean solutions that satisfy the condition (3) the Coleman instantons.
Recently, in [4, 5] a class of (unbounded) potentials in an arbitrary spacetime dimension was identified, for which the Coleman instantons do not exist. See [6, 7, 8] for new instantons which do not satisfy the Coleman boundary conditions, Eq. (3).*1*1*1The authors introduced the UV cutoff to make the action finite. In this letter we investigate another case where $O(4)$-symmetric Euclidean solutions are singular but their action remains finite.*2*2*2This situation is reminiscent of the Hawking-Turok instanton [9] where gravity is taken into account. In this case, the instanton solution diverges as $\rho\rightarrow 0$, but the action is finite. Namely, we consider $O(4)$-symmetric instantons that do not obey the second condition of Eq. (3), hence they are singular at the true-vacuum bubble. Also, similar to [4, 5], we are interested in an unbounded potential, which violates one of the assumptions for Coleman’s theorem. Moreover, we explicitly show that even if small and regular deformations are taken into account, the total action is still given by that computed on the $O(4)$-symmetric ones. We do not include gravity in this letter.*3*3*3We do not consider the backreaction on the metric due to scalar fluctuations.
@@
The above condition must be satisfied for any value of $m_{1}$ and $\Phi_{\rm M}$. Notice that in our setup we have $\Phi_{\rm M}<0$. For concreteness, in Fig. 1 we plot the potential (19) with $m_{1}=2.05\Phi_{\star}$, $m_{2}=1.45\Phi_{\star}$, $\Phi_{\rm M}=-0.54\Phi_{\star}$ and $\alpha=0.5$. Note that these values of parameters are chosen such that all the matching conditions of the background solutions are satisfied. In the plot, the red solid line corresponds to the exponential potential, whereas the blue dashed and the green dot-dashed lines refer to the potentials $V_{1}$ and $V_{2}$, respectively. The two black dots represent the matching points at $\Phi=0$ and $\Phi=\Phi_{2}=-0.50\Phi_{\star}$.
Figure 1:  $V(\Phi)/(\alpha\Phi_{\star})^{4}$ in (19) as a function of $\Phi$ with $m_{1}=2.05\Phi_{\star}$, $m_{2}=1.45\Phi_{\star}$ and $\Phi_{\rm M}=-0.54\Phi_{\star}$. The solid red line represents the exponential potential, while the blue dashed and the green dot-dashed lines refer to the potentials $V_{1}$ and $V_{2}$ in Eq. (19) respectively. The two black points in the plot refer to the two matching locations at $\Phi=0$ and at $\Phi=\Phi_{2}=-0.50\Phi_{\star}$. The false vacuum for this particular choice of parameters is located at $\Phi_{\rm M}=-0.54\Phi_{\star}$.
@@
Figure 2:  Parameter space of $m_{1}/\Phi_{\star}$ and $m_{2}/\Phi_{\star}$. The condition (28) is satisfied in the yellow region. The green line refers to the condition $\bar{\Phi}(\tilde{\rho}_{2})=\Phi_{2}$, the red curve corresponds to the condition $c_{8}=0$. We choose $\tilde{\rho}_{2}=5$ and $\alpha=0.5$. The black dot is $\{2.05,1.45\}$, for which the fluctuation is regular everywhere (their corresponding $\Phi_{2}$ and $\Phi_{\rm M}$ are respectively given by $\Phi_{2}=-0.50\Phi_{\star},\Phi_{\rm M}=-0.54\Phi_{\star}$).
@@
Figure 3:  The solutions for $\bar{\Phi}$ (solid line) and $f_{2}(\tilde{\rho})$ (dashed line). The red, blue and green colors refer to the solutions in the exponential potential, $V_{1}$ and $V_{2}$ regimes, respectively. We set $\alpha=0.5$, $m_{1}=2.05\Phi_{\star}$ and $m_{2}=1.45\Phi_{\star}$, so that $\Phi_{2}=-0.50\Phi_{\star}$ and $\Phi_{\rm M}=-0.54\Phi_{\star}$. The two matching points (black dots) are located at $\tilde{\rho}=1$ and $\tilde{\rho}_{2}=5$.
In Fig. 2 we plot the regular condition $c_{8}=0$ with $\tilde{\rho}_{2}=5$ and $\alpha=0.5$, together with the conditions (28) and $\bar{\Phi}(\tilde{\rho})=\Phi_{2}$. In the plot the red line corresponds to the vanishing of the coefficient $c_{8}$ of the solution (Beyond Coleman’s Instantons). The intersection point $\{2.05,1.45\}$ represents the location where all of the conditions [$c_{8}=0$, $\bar{\Phi}(\tilde{\rho})=\Phi_{2}$ and Eq. (28)] are simultaneously satisfied. In other words, for particular values of $\tilde{\rho}_{2}$, $a_{1}$ and $\alpha$ there exists a point where the background $\bar{\Phi}$ and the regular deformation can be realized. Therefore, our independent parameters are $\tilde{\rho}_{2}$ and $\alpha$. For other values of $\tilde{\rho}_{2}$ and $\alpha$ one needs to check the existence of the solutions case by case.
For completeness, in Fig. 3 we plot both the solutions $\bar{\Phi}$ (solid line) and $f_{2}$ (dashed line) as a function of $\tilde{\rho}$ with $\alpha=0.5$, $a_{1}=1$, $m_{1}=2.05\Phi_{\star}$ and $m_{2}=1.45\Phi_{\star}$ such that we have $\Phi_{2}=-0.50\Phi_{\star},\Phi_{\rm M}=-0.54\Phi_{\star}$. Note that in the plot we appropriately rescale $f_{2}$ with the parameter $a_{1}$. The solutions in the exponential potential, $V_{1}$ and $V_{2}$ regimes are labeled by red, blue and green respectively. The black dot denotes the matching points at $\tilde{\rho}=1$ and $\tilde{\rho}=5$. We see that $\bar{\Phi}$ is singular as $\tilde{\rho}\to 0$, while it goes to $\Phi_{\rm M}$ as $\rho\to\infty$. Moreover, the solution $f_{2}(\tilde{\rho})$ is regular everywhere, especially at the origin and infinity. Therefore, in this model there exists the regular solution for small deformations around the singular instanton.
@@
Discussion. We studied a possibility of extending the theorem [3] to the case where the $O(4)$-symmetric instanton solution exhibits a singular behavior at $\rho=0$. In particular, we considered the case in which the solution does not satisfy Coleman’s condition (3). We further investigated a small anisotropic deformation around the singular $O(4)$-symmetric instanton. We proposed a concrete example with a piecewise potential given by (19). We then analyzed the solutions for $\bar{\Phi}$, that is singular at $\rho=0$ and goes to false vacuum as $\rho\rightarrow\infty$. Finally, we studied the dynamics of small anisotropic deformation and found that there exists a small and regular deformation around the singular $O(4)$-symmetric solution, which does not contribute to the total bounce action. Our result implies that there exist non-$O(4)$-symmetric solutions with finite action beyond Coleman’s instantons, and gives rise to the possibility of the existence of a non-$O(4)$-symmetric instanton with a lower action. This opens several possibilities to study other cases, where the Coleman’s theorem may not apply. Moreover, this can be extended to the case where gravity is present (a general theorem is not known yet), see e.g. [12, 13, 14, 15]. In our separate work [10] we will discuss other types of singular behaviors in addition to logarithmic divergence, and the generic form of the instanton potential.
Acknowledgements. We thank S. Mukohyama, N. Oshita, R. Saito and K. Takahashi for useful discussions. M.S. and V.Y. are supported by World Premier International Research Center Initiative (WPI Initiative), MEXT, Japan. Y.Z is supported by the Fundamental Research Funds for the Central Universities, and by the Project 12475060 and 12047503 supported by NSFC, Project 24ZR1472400 sponsored by Natural Science Foundation of Shanghai, and also by China Scholarship Council.

### v2

@@ Abstract
where $\rho$ denotes the radial coordinate in the 4-dimensional Euclidean space. The second condition in (3) guarantees the regularity of the solution at the center of the true-vacuum bubble. We call Euclidean solutions that satisfy the condition (3) Coleman’s instantons. There have been many studies on the properties on various types of Coleman’s instantons, e.g. [4, 5, 6, 7].
Recently, a class of (unbounded) potentials in an arbitrary spacetime dimension was identified, for which the Coleman instantons do not exist [8, 9, 10]. See [11, 12, 13, 14] for new instantons*1*1*1The authors introduced the UV cutoff to make the action finite. and pseudo bounces which do not satisfy the Coleman boundary conditions, Eq. (3). In this letter we investigate another case where $O(4)$-symmetric Euclidean solutions are singular but their action remains finite.*2*2*2This situation is reminiscent of the Hawking-Turok instanton [15] where gravity is taken into account. In this case, the instanton solution diverges as $\rho\rightarrow 0$, but the action is finite. Namely, we consider $O(4)$-symmetric instantons that do not obey the second condition of Eq. (3), hence they are singular at the true-vacuum bubble. Also, similar to [8, 9], we are interested in an unbounded potential, which violates one of the assumptions for Coleman’s theorem. Moreover, we explicitly show that even if small and regular deformations are taken into account, the total action is still given by that computed on the $O(4)$-symmetric ones. We do not include gravity in this letter.*3*3*3We do not consider the backreaction on the metric due to scalar fluctuations.
@@
The above condition must be satisfied for any value of $m_{1}$ and $\Phi_{\rm M}$. Notice that in our setup we have $\Phi_{\rm M}<0$. For concreteness, in Fig. 1 we plot the potential (19) with $m_{1}=0.98\Phi_{\star}$, $m_{2}=7.11\Phi_{\star}$, $\Phi_{\rm M}=-2.35\Phi_{\star}$ and $\alpha=0.5$. Note that these values of parameters are chosen such that all the matching conditions of the background solutions are satisfied. In the plot, the red solid line corresponds to the exponential potential, whereas the blue dashed and the green dot-dashed lines refer to the potentials $V_{1}$ and $V_{2}$, respectively. The two black dots represent the matching points at $\Phi=0$ and $\Phi=\Phi_{2}=-2.34\Phi_{\star}$.
Figure 1:  $V(\Phi)/(\alpha\Phi_{\star})^{4}$ in (19) as a function of $\Phi$ with $m_{1}=0.98\Phi_{\star}$, $m_{2}=7.11\Phi_{\star}$ and $\Phi_{\rm M}=-2.35\Phi_{\star}$. The solid red line represents the exponential potential, while the blue dashed and the green dot-dashed lines refer to the potentials $V_{1}$ and $V_{2}$ in Eq. (19) respectively. The two black points in the plot refer to the two matching locations at $\Phi=0$ and at $\Phi=\Phi_{2}=-2.34\Phi_{\star}$. The false vacuum for this particular choice of parameters is located at $\Phi_{\rm M}=-2.35\Phi_{\star}$.
@@
Figure 2:  Parameter space of $m_{1}/\Phi_{\star}$ and $m_{2}/\Phi_{\star}$. The condition (28) is satisfied in the yellow region. The green line refers to the condition $\bar{\Phi}(\tilde{\rho}_{2})=\Phi_{2}$, the red curve corresponds to the condition $c_{8}=0$. We choose $\tilde{\rho}_{2}=5$ and $\alpha=0.5$. The black dot is $\{0.98,7.11\}$, for which the fluctuation is regular everywhere (their corresponding $\Phi_{2}$ and $\Phi_{\rm M}$ are respectively given by $\Phi_{2}=-2.34\Phi_{\star},\Phi_{\rm M}=-2.35\Phi_{\star}$). The small figure explicitly shows the intersection point.
@@
Figure 3:  The solutions for $\bar{\Phi}$ (solid line) and $f_{2}(\tilde{\rho})$ (dashed line). The red, blue and green colors refer to the solutions in the exponential potential, $V_{1}$ and $V_{2}$ regimes, respectively. We set $\alpha=0.5$, $m_{1}=0.98\Phi_{\star}$ and $m_{2}=7.11\Phi_{\star}$, so that $\Phi_{2}=-2.34\Phi_{\star}$ and $\Phi_{\rm M}=-2.35\Phi_{\star}$. The two matching points (black dots) are located at $\tilde{\rho}=1$ and $\tilde{\rho}_{2}=5$.
In Fig. 2 we plot the regular condition $c_{8}=0$ with $\tilde{\rho}_{2}=5$ and $\alpha=0.5$, together with the conditions (28) and $\bar{\Phi}(\tilde{\rho})=\Phi_{2}$. In the plot the red line corresponds to the vanishing of the coefficient $c_{8}$ of the solution (Beyond Coleman’s Instantons). The intersection point $\{0.98,7.11\}$ represents the location where all of the conditions [$c_{8}=0$, $\bar{\Phi}(\tilde{\rho})=\Phi_{2}$ and Eq. (28)] are simultaneously satisfied. In other words, for particular values of $\tilde{\rho}_{2}$, $a_{1}$ and $\alpha$ there exists a point where the background $\bar{\Phi}$ and the regular deformation can be realized. Therefore, our independent parameters are $\tilde{\rho}_{2}$ and $\alpha$. For other values of $\tilde{\rho}_{2}$ and $\alpha$ one needs to check the existence of the solutions case by case.
For completeness, in Fig. 3 we plot both the solutions $\bar{\Phi}$ (solid line) and $f_{2}$ (dashed line) as a function of $\tilde{\rho}$ with $\alpha=0.5$, $a_{1}=1$, $m_{1}=0.98\Phi_{\star}$ and $m_{2}=7.11\Phi_{\star}$ such that we have $\Phi_{2}=-2.34\Phi_{\star},\Phi_{\rm M}=-2.35\Phi_{\star}$. Note that in the plot we appropriately rescale $f_{2}$ with the parameter $a_{1}$. The solutions in the exponential potential, $V_{1}$ and $V_{2}$ regimes are labeled by red, blue and green respectively. The black dot denotes the matching points at $\tilde{\rho}=1$ and $\tilde{\rho}=5$. We see that $\bar{\Phi}$ is singular as $\tilde{\rho}\to 0$, while it goes to $\Phi_{\rm M}$ as $\rho\to\infty$. Moreover, the solution $f_{2}(\tilde{\rho})$ is regular everywhere, especially at the origin and infinity. Therefore, in this model there exists the regular solution for small deformations around the singular instanton.
@@
Discussion. We studied a possibility of extending the theorem [3] to the case where the $O(4)$-symmetric instanton solution exhibits a singular behavior at $\rho=0$. In particular, we considered the case in which the solution does not satisfy Coleman’s condition (3). We further investigated a small anisotropic deformation around the singular $O(4)$-symmetric instanton. We proposed a concrete example with a piecewise potential given by (19). We then analyzed the solutions for $\bar{\Phi}$, that is singular at $\rho=0$ and goes to false vacuum as $\rho\rightarrow\infty$. Finally, we studied the dynamics of small anisotropic deformation and found that there exists a small and regular deformation around the singular $O(4)$-symmetric solution, which does not contribute to the total bounce action. Our result implies that there exist non-$O(4)$-symmetric solutions with finite action beyond Coleman’s instantons, and gives rise to the possibility of the existence of a non-$O(4)$-symmetric instanton with a lower action. This opens several possibilities to study other cases, where the Coleman’s theorem may not apply. It would be interesting to explore whether the so-called pseudo bounces [10, 14] give lower or higher action than the one in our case, and study their implications.*10*10*10It could be that some off-shell configurations lead to lower actions, but they may correspond to subcritical bubbles. We leave this investigation to future work. Moreover, it is of interest to extend our analysis to the case with gravity. In connection with this, see e.g. [18, 19, 20, 21, 22, 23]. In our separate work [16] we will discuss other types of singular behaviors in addition to logarithmic divergence, and the generic form of the instanton potential.
Acknowledgements. We thank S. Mukohyama, N. Oshita, R. Saito and K. Takahashi for useful discussions. We thank J. R. Espinosa for pointing out an error in the first version of the paper. This work is supported in part by JSPS KAKENHI Nos. JP20H05853 and JP24K00624. M.S. and V.Y. are supported by World Premier International Research Center Initiative (WPI Initiative), MEXT, Japan. Y.Z. is supported by the Fundamental Research Funds for the Central Universities, and by the Project 12475060 and 12047503 supported by NSFC, Project 24ZR1472400 sponsored by Natural Science Foundation of Shanghai, and also by China Scholarship Council.